9) Which of the following imply a deflation? I. persistently increasing CPI III. positive CPI V. persistently lower inflation rate IL. persistently decreasing CPI IV, negative CPI VI. negative inflation rate C) II, IV and VI 10) The quantity theory of money argues that, in the long run (when RGDP stays constant), the percentage change in money will create an equal percentage change in A) velocity B) real GDP C) inflation rate. D) the price level. 11) If velocity is 6 and the quantity of money is $2 trillion, what is nominal GDP? A) $12 trillion B) S6 trillion C) S3 trillion D) S333 billion 12) IrV 5, P-$3, and Y- 50, then the quantity of money equals A) S10. B) S30. C) S150. D) $300 13) Suppose the money growth rate is 3 percent, velocity is constant, and price level is growing at 2 percent. What is the growth rate of real GDP? A) 1 percent B) 5 percent C) 3 percent D) 6 percent 14) Money grows faster than real GDP will result, while money grows slower than real GDP will result- A) Inflation; deflation B) Deflation; inflation D) Deflation; deflation C) Inflation; inflation 15) In the United States, the central bank is the A) Bank of America B) Federal Reserve System C) Federal Reserve Bank of New York D) Federal Reserve Bank of Washington D.C. 16) Other things constant, the quantity theory of money concludes that any increase in the quantity of money A) decreases the demand for money B) decreases in the aggregate price level. C) decreases the aggregate level of nominal income. D) proportionally increases the nominal GDP 17) The price level, which is the value that money stored, is determined by A) amount of money B) RGDP C) the value of dollar bill itself D) Both A and B 18) The value that money stored has dropped, then A) price level will increase as people need more dollar bills to buy the same good B) price level will decrease as people need less dollar bills to buy the same good C) it wont affect the price level directly D) RGDP will drop as well
Answer #1

9.

D.

When the CPI is decreasing on a regular basis, then it means that price is coming down and there is a negative inflation. So, it is a case of deflation.

10.

D

It will bring an equal percentage change in the price level.

11.

A.

Working note:

As per the quantity theory of money,

M*V = Nominal GDP

Nominal GDP = 2*6 = $12 trillion

12.

B

Working note:

As per the quantity theory of money,

M*V =P*Y

M = 3*50/5 = $30

13.

A

Growth rate in real GDP = 3% - 2% = 1%

14.

A

Relatively faster growth of money in comparison to the real GDP, will bring inflation and in the reverse case, it will bring deflation.

Pl. repost other unanswered questions for their proper answers!